This is a rational equation. 4 over X plus seven equals 2 over three X
Answer by unlockmath(1688)

Hello,
Let's write the equation out like this:
4/(x+7)= 2/3x
Let's remove the denominators on each side by multiplying (3x)(x+7)
This will leave us with:
4(3x) = 2(x + 7)
or rewritten 12x = 2x + 14
Move 2x by subtracting both sides
10x = 14
Divide each side by 10
x = 14/10 reduce it to 7/5 or 1 2/5
RJ Toftness
www.math-unlock.com